Michael likes to go for daily jogs with his dog, Max. If the weather is nice, he is 90% likely to jog for 5km. If the weather is rainy, he is only 55% likely to jog for 5km. The weather forecast for tomorrow indicates 40% chance of rain. Determine the probability that Michael will jog for 5km.


1
Expert's answer
2021-01-14T17:42:30-0500

Let "E_1,E_2" and "A" are three events defined as follows:

"E_1=" Weather forecast is nice

"E_2=" Weather forecast is rainy

"A=" Michael will jog for "5" km.

Then "P(E_1)=\frac{60}{100}"

"P(E_2)=\frac{40}{100}"

and "P({A}/{E_1})= \frac{90}{100}"

"P({A}/{E_2})= \frac{55}{100}"

Then the required probability that Michael will jog for "5" km is

"P(A)=P(E_1).P({A}\/{E_1})+P(E_2).P({A}\/{E_2})"

"=\frac{60}{100}.\frac{90}{100}+\frac{40}{100}.\frac{55}{100}"

"=\frac{54}{100}+\frac{22}{100}"

"=\frac{76}{100}"

"=0.76"